# Discovering Effective Policies for Land-Use Planning

### 13 Dec 2023 v2

@@ 5.1 Prediction
In preliminary experiments, it turned out to be difficult to achieve high accuracy globally in a single model. Therefore, separate models were trained on different subsets of countries: Western Europe, South America, and the United States. The Western Europe subset contains the UK, France, Germany, the Netherlands, Belgium, Switzerland, and Ireland. The South American subset includes Brazil, Bolivia, Paraguay, Peru, Ecuador, Colombia, Venezuela, Guyana, Suriname, Uruguay, Argentina, and Chile. These countries were chosen as a representative subset to make the experiments computationally less expensive.
@@
Upon examining the dataset, a linear relationship was found between ||||ELUC|||| and change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use and cell area were removed from the inputs. Thus, the differences in land use were used as the only input features in the experiments. Data from years [1851-2011] was used to train the LinReg models.
Second, a random forest model (RF) was trained on the same data, with the default scikit-learn parameters and the full set of input features. To make training time feasible, the models for Western Europe, South America, and the United States were trained on data from the years [1982-2011] and the global models were trained with data from [2010-2011]. Preliminary experiments showed that constraining the training set to more modern years did not significantly decrease performance.
@@
Third, a neural network (neural net) was trained for the same task. For Western Europe, South America, and the United States, it was trained on the same data as LinReg; For global, the neural net was trained on a random 10% of the global data, due to training time limitations. The neural nets were fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer that took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016).
After training, all models were tested with data from [2012-2021]. As shown in Table 1, the LinReg models performed consistently the worst, suggesting that the problem includes significant nonlinear dependencies after all. RF performed significantly better; however, RF models do not extrapolate well beyond the training examples, which turns out to be important when evolving prescriptors for extreme land-use changes. In contrast, neural nets both capture nonlinearities and extrapolate well, and turned out to be best models overall. In each region the neural network trained on that region performed the best, except in SA, where it was close. The LinReg and neural network have many fewer parameters than RF (taking 4KB and 440KB on disk respectively vs. 2GB for RF), and LinReg trains the fastest.
@@
Figure 2: Visualizing the differences in model behavior. Predictions are created for the global models using synthetic data that consists of changing 100% of land type A (row) to 100% of type B (column). The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to extreme values, resulting in low predictions; LinReg and Neural Net are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
@@
In order to visualize the behavior of the predictor models, they were queried by every possible conversion from 100% type A to 100% type B (the types primf, primn, and urban were excluded from B and c4per from A and B for reasons outlined in Section 5.2). Figure 2 shows the result for the global models. The main result is that all models generally rank crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and Neural Net make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to extrapolate well to extreme changes in land usage, resulting in low-magnitude ELUC values.
@@
Thus, since the neural network achieves better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the global neural network was used as the surrogate to evolve the prescriptors.
5.2 Prescriptions
@@
In the experiments in this paper, the prescriptors were fully connected neural networks with two layers of weights, implemented with Keras/Tensorflow. The topology was fixed and only the weights were evolved. The input consisted of 13 float values, i.e. the 12 land-usage percentages and the area of the geographical cell. The hidden layer contained 16 units with tanh activation.
@@
During evolution, prescriptor candidates were evaluated on a random subset of 1% of the global dataset from the years [1851-2011] with the global neural net predictor. Prescriptor output was a single vector of size eight (the 12 land usage percentages minus primf, primn, and urban, which were considered fixed, and c4per which was found to occur too rarely in the dataset to predict reliably), and a softmax activation. These eight outputs specify the prescriptor’s suggestion for the percentage of land usage of each type. The eight outputs were combined with the four fixed land-usage values and the cell area and passed to the Predictor model to get the predicted ‘ELUC’ metric. This metric is aggregated across the evaluation data set to form the performance outcome of the suggestion. The difference between the eight outputs and the corresponding inputs is used to compute the percentage of land-use change, i.e. the cost outcome of the suggestion.
@@
The aggregated ELUC and the change outcomes were minimized with an evolutionary search on the neural net weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II (Deb et al., 2000) to account for multiobjectivity. To speed up evolution, two prescriptor models were trained with backpropagation: One to prescribe no changes and the other to prescribe as much secdf as possible (which was observed to reduce ELUC the most); these two models were injected into the initial population. The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of 0 and a standard deviation of 1 (Saxe et al., 2014). Evolution was run for 100 generations with the following parameters:
@@
Figure 3: Evolution of prescriptors with the global neural network predictor. (Left) The Pareto front moves towards the lower left corner over evolution, finding better implementations for the different tradeoffs of the ELUC and change objectives. (Right) Each prescriptor evaluated during evolution is shown as a dot, demonstrating a wide variety of solutions and tradeoffs. The final Pareto front is shown as red dots in both figures, constituting a set of solutions from which the decision-maker can choose a preferred one.
@@
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and change objectives. They represent a wide variety of tradeoffs, and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to be implemented.
Figure 4: The Pareto fronts of evolved prescriptors vs. heuristic baselines, with ELUC and change evaluated on the global test set. The trained prescriptors achieve better solutions than the baselines in the middle-change region where the changes matter the most, demonstrating that they can take advantage of nonlinear relationships in land use to discover useful, non-obvious solutions.
@@
To demonstrate the value of evolutionary search, the prescriptors in the Pareto front were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple “even” heuristic, where an equal proportion of land equal to the threshold is taken from all land types and converted to secdf. The second is the optimal linear heuristic, where as much land as possible is first taken from the land type that has the highest weight in the global LinReg model, then from the land type that has the second highest weight, and so on until the given change threshold is reached. Ten such heuristic prescriptors were created with evenly distributed change thresholds in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the global test set from [2012-2021].
@@
As Figure 3 shows, all prescriptions converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the evolved prescriptors outperformed both heuristics, thus demonstrating that they are exploiting the nonlinear relationships well.
5.3 Interactive Evaluation
An interactive demo of the trained models is provided at https://landuse.evolution.ml. It allows the user to explore different locations, observe the prescribed actions for them, and see their outcomes. It is also possible to modify those actions and see their effect, thus evaluating the user’s skills as a decision-maker compared to that of machine learning.
As an example, Figure 5 shows the grid cells from which the user has selected one. The system’s recommends a change in land-use that results in a 26.5% reduction in carbon emissions with a 22.0% change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
@@ 5.3 Interactive Evaluation
Figure 5: A suggested land-use change for a given location. ($a$) The location is indicated by the red dot among the UK grid cells. ($b$) One prescriptor is chosen from the Pareto front spanning minimal change and minimal ELUC (top left). The current land use is shown on the left pie chart and the recommended one on the right pie chart. This prescriptor recommends decreasing pasture and annual C3 crops and increasing secondary forest and secondary non-forest vegetation, resulting in a 26.5% decrease in carbon emissions with a 22.0% change. The user can then select different solutions from the Pareto front and modify the sliders manually to explore alternatives.
6 Future Work
@@
The most immediate future work consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be benefit of including LinReg, RF, and neural network versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. Also, a metric will be developed for prescriptor performance, such as the hypervolume of the Pareto front, making it possible to compare different predictor and prescriptor approaches quantitatively.
@@
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production. Similarly, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
7 Conclusion
Land-use policy is an area of climate change where local decision-makers can have a large impact. In this paper, historical data and simulation technology are brought together to build an efficient machine-learning model that can predict the outcomes of these decisions efficiently for different contexts. An evolutionary optimization process is then used to identify effective policies, balancing the cost of land-use change and its effect in reducing carbon emissions. Machine learning methods can thus play an important role in empowering decision-makers on climate change issues.

### 09 Feb 2024 v3

@@ 5.1. Prediction
In preliminary experiments, prediction performance was found to differ between major geographical regions. To make these differences explicit, separate models were trained on different subsets of countries: Western Europe (EU), South America (SA), and the United States (US), making up 1.12, 9.99, and 7.26 percent of the global dataset, respectively. The EU subset contains the UK, France, Germany, the Netherlands, Belgium, Switzerland, and Ireland. The SA subset includes Brazil, Bolivia, Paraguay, Peru, Ecuador, Colombia, Venezuela, Guyana, Suriname, Uruguay, Argentina, and Chile. These countries were chosen as a representative subset to make the experiments computationally less expensive.
@@
Upon examining the global dataset, a linear relationship was found between ||||ELUC|||| and land-use change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use, latitude, longitude, cell area, and time were removed from the inputs. Thus, the difference in land use was used as the only input feature in the experiments. Data from the years 1851–2011 (inclusive), totalling 39,209,670 samples globally, was used to train the LinReg models; the 2,420,350 remaining samples from 2012–2021 (inclusive) were used for testing.
Second, a random forest model (RF) was trained on the same data but now including the full set of input features, in order to take advantage of RF’s ability to decide which features are most useful. The forest consisted of 100 decision trees with unrestricted maximum depth; at each split, a random subset of $n$ features was considered (where $n$ is the square root of the total number of features). To make the training time of the Global RF model feasible, it was trained with data from only 1982–2011. Preliminary experiments showed that constraining the training set in this manner did not decrease performance significantly. Both the linear regression and random forest models were trained using the scikit-learn library Pedregosa et al. (2011).
@@
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library Paszke et al. (2019). The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016). It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm Loshchilov and Hutter (2019). After each epoch the learning rate was scaled by a factor of 0.1.
Refer to caption ($a$) LinReg predictions
Refer to caption ($b$) RF predictions
Refer to caption ($c$) NeuralNet predictions
@@
Figure 2. Visualizing the differences in model behavior. Predictions are created for the Global models using synthetic data created by changing 100% of land-use type A (row) to 100% of type B (column) in a 1% sample across the range of latitude, longitude, cell area, and year occurring in the training data. The results are averaged for each conversion from A to B. The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to large values, resulting in low predictions (shown in a much smaller color scale to still make them visible); LinReg and NeuralNet are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg and NeuralNet in all regions except Western Europe. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions, as well as RF in Western Europe. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
@@
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the predictor takes latitude, longitude, cell area, and year as input, 392,096 (1%) latitude/longitude/cell-area/year combinations from the training dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes, and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
@@
Thus, since NeuralNet achieves better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
5.2. Prescriptions
@@
The prescriptors were fully connected neural networks with two layers of weights, implemented with Keras/Tensorflow. The topology was fixed and only the weights were evolved. The input consisted of 12 float values, i.e. the eight land-use percentages and the latitude, longitude, and area of the geographical cell, as well as the year. The hidden layer contained 16 units with tanh activation.
Refer to caption ($a$) Evolution of the Pareto front
Refer to caption ($b$) Performance of all evaluated prescriptors
Figure 3. Evolution of prescriptors with the Global NeuralNet predictor. ($a$) The Pareto front moves towards the lower left corner over evolution, finding better implementations for the different tradeoffs of the ELUC and land-use change objectives. ($b$) Each prescriptor evaluated during evolution is shown as a dot, demonstrating a wide variety of solutions and tradeoffs. The final Pareto front (collected from all generations) is shown as red dots, constituting a set of solutions from which the decision-maker can choose a preferred one.
Figure 4. The Pareto fronts of Evolved Prescriptors vs. heuristic baselines, with ELUC and land-use change evaluated on the Global test set. The Evolved Prescriptors achieve better solutions than the baselines in the middle-change region where the land-use changes matter the most, demonstrating that they can take advantage of nonlinear relationships in land use to discover useful, non-obvious solutions.
Refer to caption ($a$) Average performance
Refer to caption ($b$) Performance on underlying samples
Refer to caption ($c$) Performance differences
Figure 5. Comparing Evolved Prescriptors with the heuristics. ($a$) The average performance of the Evolved Prescriptor dominates those of the heuristics. ($b$) The averages are expanded into actual samples in the test set (subsampled for readability). The Evolved Prescriptor suggests many more large changes than the heuristics. ($c$) The differences in change percentage and ELUC between the Evolved Prescriptor and the Perfect Heuristic for each test sample, with color indicating which solution dominates. Surprisingly, the Evolved Prescriptor dominates the Perfect Heuristic only on a few individual samples. Thus, evolution discovered the insight that it is possible to do well globally by utilizing a few cases where large change is possible.
Figure 6. Characterizing the recommendations. The strategies used by the Evolved Prescriptors and the heuristics are illustrated by plotting the average amount of change in each land-use type. Primarily, they all suggest converting cropland to secondary forest. The Evolved Prescriptor removes more crop on average, and generally takes advantage of more change.
@@
During evolution, prescriptor candidates were evaluated on a random subset of 1% of the Global dataset from years 1851–2011 with the Global NeuralNet predictor. Prescriptor output was a single vector of size five (the eight land-use percentages minus primf, primn, and urban, which were considered fixed), produced by the ReLU activation function. These five outputs were normalized to sum to the previous amount of land taken by these five types. The result is the prescriptor’s suggestion for the percentage of land use of each type. The five outputs were combined with the fixed land-use values for primf, primn, and urban to compute the difference from the initial land use. This difference was concatenated with the initial land use as well as latitude, longitude, cell area, and year, and passed to the Predictor model to get the predicted ELUC metric. This metric was aggregated across the evaluation data set to obtain the performance outcome of the suggestion. The cost outcome, i.e. the percentage of land-use change, was computed based on the difference between the five outputs and the corresponding inputs. All positive differences were summed for a given cell and divided by the total land-use percentage of the cell (excluding nonland).
@@
The aggregated ELUC and the land-use change outcomes were minimized with an evolutionary search on the neural network weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II (Deb et al., 2000) to account for multiobjectivity. To speed up evolution, two prescriptor models were trained with backpropagation and injected into the initial population: One to prescribe no land-use changes at all and the other to prescribe as much secondary forest (secdf) as possible (which was observed to reduce ELUC the most). The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of zero and a standard deviation of one (Saxe et al., 2014). Evolution was run for 100 generations with the following parameters:
@@
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and land-use change objectives. These solutions will be referred to as "Evolved Prescriptors" in the rest of the paper. They represent a wide variety of tradeoffs and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to implement.
@@
To demonstrate the value of evolutionary search, the Evolved Prescriptors were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple baseline called "Even Heuristic", where an equal proportion of land up to the threshold is taken from all land-use types and converted to secdf. The second is an optimal linear baseline called "Perfect Heuristic", where as much land as possible is first taken from the land-use type that has the highest weight in the Global LinReg model, then from the land-use type that has the second highest weight, and so on until either the given land-use change threshold is reached or there is no land left to convert to secdf. One hundred such heuristic solutions were created with evenly distributed land-use change thresholds between 0% and 100% in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the Global test set from the years 2012–2021.
@@
As Figure 3 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, thus demonstrating that they are likely exploiting the nonlinear relationships well. The performance for each prescription method can be quantified in terms of hypervolume, which is measured as the rectangular area above and to the right of each point (only counting overlapping area once), extending to ELUC = 0 and change = 1. The Even Heuristic is outperformed by the Perfect Heuristic which is outperformed by the Evolved Prescriptors, with hypervolumes of 18.24, 18.87, and 19.13, respectively.
5.3. Prescriptor Behaviors
[…]
Refer to caption ($a$) The location of the cell
Refer to caption ($b$) Suggestions of one Evolved Prescriptor chosen from the Pareto front
@@ 5.3. Prescriptor Behaviors
Figure 7. A suggested land-use change for a given location (screenshots from the demo at https://landuse.evolution.ml). ($a$) The location is indicated by the red dot among the UK grid cells. ($b$) One Evolved Prescriptor is chosen from the Pareto front spanning minimal change and minimal ELUC (top left). The current land use is shown on the left pie chart and the recommended one on the right pie chart, as well as the sliders on the left. This prescriptor recommends decreasing pasture and crops and increasing secondary forest, resulting in an 18.09 tC/ha decrease in carbon emissions with a 15.77% land-use change. The user can then select different solutions from the Pareto front and modify the sliders manually to explore alternatives.
[…]
The reason is that evolution discovered a way to take advantage of the large changes. In order to perform better on average, it does not need to outperform the Perfect Heuristic at the individual cell level; instead, it can identify and utilize cases where large changes make a big difference. Thus, even though the Evolved Prescriptor almost never beats the Perfect Heuristic in any individual cell, when averaged across all cells, it dominates the heuristic. In other words, evolution discovered the insight that it is most important to focus land-use changes in those regions where the largest changes are possible. Pick your battles right, and you’ll win overall.
In order to see what the recommendations look like, the average policies of the three approaches are illustrated in Figure 6. The greatest amount of negative change made is to crop and positive is to secondary forest, i.e. all approaches recommend converting cropland to forest, which makes sense.
5.4. Interactive Evaluation
An interactive demo of the trained/evolved models is provided at https://landuse.evolution.ml. It allows the user to explore different locations, observe the prescribed actions for them, and see their outcomes. It is also possible to modify those actions and see their effect, thus evaluating the user’s skills as a decision-maker compared to that of machine learning.
As an example, Figure 7 shows the grid cells from which the user has selected one. The system recommends a change in land-use that results in an 18.09 tC/ha reduction in carbon emissions with a 15.77% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
6. Future Work
In the project so far we have not distinguished between the different types of crops and only partly considered regional variations in how land-use changes affect carbon stocks in soil and vegetation (by training different models for different regions). The assumed values for carbon stocks in soil and vegetation represent conditions of the recent past (Hansis et al., 2015); they do not consider environmental effects such as disturbances due to fires or droughts or better vegetation growth due to CO2 effects or a warmer climate, which can substantially impact terrestrial CO2 fluxes (e.g. Aragão et al., 2018; Pongratz et al., 2021; Rosan et al., 2024). Further, estimates of CO2 fluxes from land-use change also depend on the spatial resolution of the land-use change dataset (Ganzenmüller et al., 2022; Bultan et al., 2022). Extending the project with such details should lead to more precise and more fine-scale estimates of the impact of land-use change decisions in future work.
@@ 6. Future Work
Another immediate future work direction consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be benefit of including LinReg, RF, and NeuralNet versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. The hypervolume of the Pareto front can be used to compare the different approaches quantitatively.
@@
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land-use type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production. Similarly, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
7. Conclusion
Land-use policy is an area relevant to climate change where local decision-makers can have a large impact. In this paper, historical data and simulation technology are brought together to build an efficient machine-learning model that can predict the outcomes of these decisions efficiently for different contexts. An evolutionary optimization process is then used to identify effective policies, balancing the cost of land-use change and its effect in reducing carbon emissions. Machine learning methods can thus play an important role in empowering decision-makers to act on climate change issues.